# Searching for periodic signals and quasi-periodic oscillations from an extremely active cycle of FRB 20240114A

### v1

@@ 2.1.2 Long-Timescale Periodicity Search
The long-timescale search aims to detect periodicity on timescales of the order of days, which corresponds to the active period of the FRB. Two complementary techniques are employed in this study: epoch folding with $\chi^{2}$ testing and Lomb-Scargle periodogram (LSP) analysis.
Figure 2: Folding results of the corresponding TOAs, based on the parameters in Table 1. From top to bottom: profiles of the three candidate signals and their corresponding time-phase plots.
The epoch folding search spans $P_{\rm min}=1.5$ d to $P_{\rm max}=T/3$ d, where $T=214.46$ d is the time difference between the first and last TOA. The frequency step size is $df=0.1/T$, which determines the corresponding period step size through the relation between $P$ and $f$. We compute weighted mean TOA for each observation, with weights proportional to the daily burst rate. The $\chi^{2}$ statistic evaluates the significance of periodicity:
$\chi^{2}=\sum_{i=1}^{n}\frac{(N_{i}-\langle N\rangle_{i})^{2}}{\langle N\rangle_{i}},$ (6)
where $N_{i}$ represents the weighted TOA counts in phase bin $i$, $\langle N\rangle_{i}=aT_{i}$ is the expected count under the null hypothesis of constant rate, with $T_{i}$ being the exposure time and $a$ the mean burst rate. In our periodicity analysis, we systematically tested phase bin configurations ranging from 4 to 16 bins. The candidate selection process incorporated several rigorous criteria: the first of which was that only signals exceeding the $3\sigma$ confidence level (with appropriate multiple testing corrections) were to be considered significant. When detecting multiple adjacent period candidates above the threshold, we merged these and retained solely the most statistically significant period. To ensure physical validity, the candidates were required to appear consistently in at least two independent bin configurations. Furthermore, we implemented a sidereal day rejection filter to exclude any spurious signals matching integer multiples ($N=1$-$10$) of the characteristic 0.99727-day sidereal period within a specified tolerance window.
Figure 3: Long-timescale periodicity search results for the TOA data. Panel a: Epoch folding search result, with green dash-dotted and red dashed lines representing the $3\sigma$ and $5\sigma$ thresholds, respectively. The blue solid line indicates the degrees of freedom. The red solid line denotes a signal exceeding the $3\sigma$ threshold around 69.18 days. Panel b: Similar to panel a, with the difference that it uses the Lomb-Scargle algorithm, detecting a $3\sigma$ excess signal around 67.51 days. Panel c: Profile obtained by folding the TOA data with a period of 69.18 days. The black bar and red dashed line indicate the profile with and without exposure time correction, respectively. Panel d: The relative exposure time distribution as a function of phase, calculated for the signal’s period.
For the LSP analysis, we use the weight of each observed TOA as the intensity and the average TOA from each observation as the time. We then conduct a periodicity search within the same range mentioned above. We use the False Alarm Probability (FAP) to quantify the detection significance. The FAP is defined as (VanderPlas, 2018)
@@
where $P_{\rm single}(z)=(1-z)^{\frac{N-1}{2}}$ represents the cumulative probability of observing a peak less than z at a specific frequency, assuming Gaussian noise. $N_{\rm eff}=f_{\rm max}T$, where $T$ is the time span of the time series, and $f_{\rm max}$ is the maximum frequency of the search.
@@ 2.2.1 Short-Timescale Periodicity
We first performed a joint search of all observational data, which did not yield any candidates meeting our selection criteria. Subsequently, we conducted an analysis of individual observations by grouping those taken within 2–3-day intervals. This clustering approach resulted in 14 distinct groups of TOA data being analyzed, none of which produced candidates that satisfied our criteria. The individual observation analysis revealed several candidates that met our selection criteria. We require each observation to include at least three TOAs to be included in the search, which resulted in the exclusion of one observation from MJD 60344 out of the total 57 observational datasets. We detected 2765 candidates exceeding the $3\sigma$ significance level after multiple testing correction. Given the wide search parameter space and sparse TOA distribution, such a number of apparent candidates is statistically expected. Therefore, we imposed stringent criteria to isolate only the most robust signals: cross-epoch recurrence and minimum $P$-value within frequency clusters. First, for candidates with similar $f$ (where similarity is defined as $\Delta f/f<10^{-3}$) yet different $\dot{f}$, only the candidate with the smallest $P$-value was retained. Subsequently, we retained candidates that reappeared in at least two separate observation epochs. This rigorous process ultimately identified three candidates exhibiting cross-epoch recurrence. The dates on which these signals appeared are marked in the same color in Figure 1.
For precise parameter determination, we performed a refined search in their $f$-$\dot{f}$ parameter space using the $Z_{n}^{2}$ statistic. Specifically, for each candidate signal, we searched for 16,000 combinations of $f$ and $\dot{f}$ centered on the values determined by the TDA, with step sizes chosen such that the maximum phase drift over the entire continuous observation period would be less than 0.01. The $Z_{n}^{2}$ analysis led to the identification of three distinct candidates: one with a period of $\sim$0.673 s (detected in observations 60401 and 60403), another with a period of $\sim$0.635 s (detected in observations 60398 and 60519), and a third with a period of $\sim$0.536 s (detected in observations 60417 and 60517). Table 1 presents the sifted candidates identified by the TDA through rigorous selection criteria, alongside their refined parameters obtained via precision analysis with the $Z_{n}^{2}$ statistic. Figure 2 displays the profile and the corresponding time-phase diagram using these optimized parameters.
@@ 3.1 Method
For periodicity detection of bursts, we adopted the Bayesian framework proposed by Huppenkothen et al. (2013). The method begins by computing the periodogram of burst time series via the FFT, then fits a null hypothesis model to obtain a maximum a posteriori (MAP) parameters. We calculate residuals $R_{j}=2I_{j}/S_{j}$, where $I_{j}$ represents the observed periodogram and $S_{j}$ denotes the null model’s prediction. Markov Chain Monte Carlo (MCMC) sampling generates simulated periodograms, establishing a distribution of maximum residuals against which observed residuals are compared for significance assessment. In order to detect QPOs, we further introduce a nested alternative hypothesis model, computing the likelihood ratio test (LRT) statistic $T_{\mathrm{LRT}}^{\mathrm{obs}}$ and determining its significance through simulated periodograms.
@@
For periodicity detection of burst clusters, we employed the ACF method proposed by Kramer et al. (2024). The method first requires determining an appropriate time window. By analyzing the distribution of burst waiting times, we found that the distribution exhibits a bimodal structure with a transition around 0.2 seconds. Based on this, bursts with intervals shorter than 0.2 seconds are clustered into the same time series. Periodic signals in ACF analysis manifest as equally spaced local maxima. We quantify the strength of periodicity using the statistic $\mathcal{P}=\sum\rho(P_{\mu}\cdot i)$, where $\rho$ denotes the power of the ACF, and $P_{\mu}\cdot i$ represents integer multiples of the period $P_{\mu}$. The original subpulse counts, widths, and amplitudes are preserved, while uniformly distributed perturbations are incorporated at periodic intervals. The simulated sequences then undergo the same ACF analysis to generate a distribution for comparison with the observed values, thereby obtaining the $P$-value.
@@
The periodicity detection method for individual bursts circumvents the challenges of decomposing complex signals into subcomponents. Conversely, the method for burst clusters leverages the timing characteristics of discrete events to facilitate the assessment of periodic significance. For details of the methodology, see references Chime/Frb Collaboration et al. (2022); Kramer et al. (2024) for the ACF-based analysis and Appendix A for the implementation of the FFT-based approach.
Figure 5: Example of a QPO embedded in a burst time series (No. 1 in Table LABEL:tab:QPO_pars). A1: Periodogram of the time series before DM removal. A2: Power spectrum of the time series after DM removal with a DM of 528.5 $\rm pc/cm^{3}$. Two peaks can be clearly seen around 260Hz and 500Hz. The fitting results using three Gaussian models (one fitting the low-frequency broadband component and two fitting the QPO) and a constant model (fitting white noise) are shown. A3: Variation curve of the relative intensity of the QPO with a central frequency of around 260Hz with respect to different DMs. B1: Dynamic spectrum of the burst, with red short lines indicating the frequency channels where RFI has been removed. B2: Fitting results of the burst time series using an asymmetric Gaussian model and a constant model. B3: Residuals of the burst time series fit. C1: The QPO parameters obtained from the fit in A2. C2: Periodogram of the burst time series after the best fit in B2, showing that the QPO is still present. C3: Periodogram of the residuals shown in B3, with no significant QPO observed. Based on C2 and C3, it is clear that the QPO signal exists in the pulse profile, rather than in the low-amplitude oscillations of the residuals.
@@ 3.2 Results
For the ACF analysis, we clustered bursts using a 0.2-second time window, resulting in 516 time series. The ACF was calculated for each of these time series, and the minimum lag of the significant peaks was identified as the candidate period $P_{\mu}$. The significance of these periodicities was assessed using the methodology outlined in Section 3.1. We identified 4 candidates with significance greater than $3\sigma$. These burst time series, their dynamic spectra, ACFs, and distributions of simulated statistics are shown in Figure 4. These time series exhibit relatively few peaks, with those peaks spaced at regular intervals. Most of these candidates exhibit 3 peaks, with one candidate showing 4 peaks, resulting in a high significance of up to $3.9\sigma$. In Appendix C (Figure 10), we also list 7 additional candidates with significance lower than $3\sigma$ yet still showing some evidence of periodicity. Notably, some of these candidates match the period well up to 3 peaks, although the agreement breaks down after the 4th peak, leading to lower significance. This suggests that the candidates with significance greater than $3\sigma$ may not exhibit periodic patterns in longer time series. The periods of these candidates range from 5 ms to 62 ms, reflecting the clustering of bursts within a 0.2-second time window. As expected, increasing the time window length makes it more difficult to maintain periodicity. This aligns with the results from searching for short-timescale periods in the TOA data, where pulse phases exhibited considerable spread rather than remaining confined to a narrow range. The 0.2-second clustering assumption implies that bursts occurring within such short intervals are likely to be physically connected, indicating the potential for periodicity in these local bursts.
Figure 6: Posterior distributions of the null hypothesis model parameters and the likelihood ratio test. Upper panel: The corner plot displays the posterior distributions of the model parameters $\mu$, $\sigma$, and $c$ from the null hypothesis employed to fit the low-frequency component and white noise in the periodogram (see Figure 5A2). Here, $\mu$, $\sigma$, and $c$ represent the amplitude, centre value, width of the Gaussian function, and white noise level, respectively. Lower panel: A comparison of the observed likelihood ratio with the distribution of $2\times 10^{6}$ likelihood ratios obtained from the posterior distributions. The observed value significantly deviates from the sample distribution, indicating a rejection of the null hypothesis. For further details, refer to Section 3.2.
[…]

### v2

@@ 2.1.2 Long-Timescale Periodicity Search
The long-timescale search aims to detect periodicity on timescales of the order of days, corresponding to the activity cycle of the FRB. The primary analysis employs Lomb-Scargle periodogram (LSP) analysis, with epoch folding used for verification of candidate signals.
For the long-timescale search, we include all TOAs from observations between 2024-01-28 and 2025-10-01 (UT) to maximize the time span and improve sensitivity to multi-month periods. The extended dataset provides reliable TOAs and burst rates suitable for long-timescale analyses. This approach ensures both high data quality and optimal sensitivity for detecting multi-month periodicity. The complete FAST observational log used for this analysis is provided in B.
For the LSP analysis, the weighted mean TOAs from each observation are used as the time series, with weights proportional to the daily burst rate. The search is conducted over frequencies corresponding to periods from $P_{\rm min}=1.5$ days to $P_{\rm max}=T_{\rm span}/2\approx 306.16$ days, where $T_{\rm span}=612.33$ days is the total time span of the observations. The frequency grid is sampled with 100 points per independent peak, resulting in a frequency step of
$df=\frac{f_{\rm max}-f_{\rm min}}{N_{\rm samp}\cdot N_{\rm peak}},$ (6)
where $f_{\rm min}=1/P_{\rm max}$, $f_{\rm max}=1/P_{\rm min}$, $N_{\rm samp}=100$ is the number of samples per peak, and $N_{\rm peak}$ is the estimated number of independent peaks within the frequency range. This ensures that narrow peaks can be resolved while covering the full period range of interest. The LSP power is computed using the standard Lomb-Scargle formulation. The significance of detected peaks is quantified using the false alarm probability (FAP) ([2018ApJS..236...16V]):
@@
where
$P_{\rm single}(z)=(1-z)^{\frac{N-1}{2}},$ (8)
NN is the number of data points, and $N_{\rm eff}=f_{\rm max}T_{\rm span}$ is the effective number of independent frequencies.
Candidate periods are selected based on the FAP significance level, with a threshold corresponding to $3\sigma$ confidence (FAP <0.0027<0.0027). For verification of these LSP-detected candidates, we employ epoch folding to examine the phase profile.
Folded profiles can be modeled with one or more Gaussian functions plus a constant baseline:
$f(x)=\sum_{k=1}^{N}A_{k}\exp\left[-\frac{(x-\mu_{k})^{2}}{2\sigma_{k}^{2}}\right]+C,$ (9)
where $N$ is the number of Gaussian components, $A_{k}$ the amplitude of the $k$-th component, $\mu_{k}$ its phase center, $\sigma_{k}$ its width, and $C$ the constant baseline. Exposure correction is applied when necessary, and the profile may be repeated or smoothed to facilitate visualization and comparison.
Figure 3: Long-timescale periodicity search results for FRB 20240114A. The analysis uses all TOAs from observations between 2024-01-28 and 2025-10-01 ($T_{\rm span}=612.33$ days), searching periods from 1.5 days to $T_{\rm span}/2\approx 306.16$ days. Candidate periods were selected at $3\sigma$ confidence (FAP <0.0027<0.0027) and verified through epoch folding with Gaussian profile modeling. A: Burst rate evolution with time, showing the count rate (counts/hour) versus MJD. The green-shaded regions mark the predicted enhanced active phases. These are derived from a phase-folding analysis using the 143.40-day period, with a Gaussian fit (panel C1) applied to identify the timing and duration of the expected enhanced activity windows. B: Lomb-Scargle periodogram showing normalized power versus period. The black solid line represents the burst rate-weighted analysis, while the orange dashed line shows the observing window function. Two significant periods are identified: Period 1 = $143.40\pm 7.19$ days (red dashed line, $5.2\sigma$ significance) and Period 2 = $73.60\pm 2.45$ days (green dashed line, $3.3\sigma$ significance). C1-C3: Phase-folded analysis for Period 1. C1: Intensity profile showing exposure-corrected (green) and uncorrected (blue) rates with Gaussian fit (red dashed), from which the enhanced active window timing and duration are determined. C2: Relative exposure time distribution across phase bins. C3: TOA versus phase, with point sizes scaled by burst rate weights. D1-D3: Similar to C1-C3 but for Period 2. Based on the analysis results, Period 2 is identified as a harmonic of Period 1.
@@ 2.2.1 Short-Timescale Periodicity
We first performed a joint search of all observational data, which did not yield any candidates meeting our selection criteria. Subsequently, we conducted an analysis of individual observations by grouping those taken within 2–3-day intervals. This clustering approach resulted in 14 distinct groups of TOA data being analyzed, none of which produced candidates that satisfied our criteria. The individual observation analysis revealed several candidates that met our selection criteria. We require each observation to include at least three TOAs to be included in the search, which resulted in the exclusion of one observation from MJD 60344 out of the total 57 observational datasets. We detected 2765 candidates exceeding the $3\sigma$ significance level after multiple testing correction. Given the wide search parameter space and sparse TOA distribution, such a number of apparent candidates is statistically expected. Therefore, we imposed stringent criteria to isolate the most reliable signals: cross-epoch recurrence and minimum $P$-value within frequency clusters. First, for candidates with similar $f$ (where similarity is defined as $\Delta f/f<10^{-3}$) but different $\dot{f}$, only the candidate with the smallest $P$-value was retained. Subsequently, we retained candidates that reappeared in at least two separate observation epochs. This rigorous process ultimately identified three candidates exhibiting cross-epoch recurrence. The dates on which these signals appeared are marked in the same color in Figure 1.
For precise parameter determination, we performed a refined search in their $f˙\dot{f}$ parameter space using the $Z_{n}^{2}$ statistic. Specifically, for each candidate signal, we searched for 16,000 combinations of $f$ and $\dot{f}$ centered on the values determined by the TDA, with step sizes chosen such that the maximum phase drift over the entire continuous observation period would be less than 0.01. The $Z_{n}^{2}$ analysis led to the identification of three distinct candidates: one with a period of $\sim$0.673 s (detected in observations 60401 and 60403), another with a period of $\sim$0.635 s (detected in observations 60398 and 60519), and a third with a period of $\sim$0.536 s (detected in observations 60417 and 60517). Table 1 presents the sifted candidates identified by the TDA through rigorous selection criteria, alongside their refined parameters obtained via precision analysis with the $Z_{n}^{2}$ statistic. Figure 2 displays the profile and the corresponding time-phase diagram using these optimized parameters.
@@ 2.2.2 Long-Timescale Periodicity
After conducting the periodicity search on the data from 2024-01-28 to 2025-10-01 (UT), we obtained two periodic candidates. Figure 3A shows the burst rate evolution with time, displaying the count rate (counts/hour) versus MJD. The Lomb-Scargle periodogram (Figure 3B) identified two prominent peaks exceeding the $3\sigma$ confidence threshold (FAP << 0.0027). The burst rate-weighted periodogram (black solid line) shows clear detections at both periods, while the window function (orange dashed line) confirms these are not artifacts of the observational sampling.
The primary period, designated as Period 1, was detected at $143.40\pm 7.19$ days with a significance of $5.2\sigma$. The secondary period, Period 2, was found at $73.60\pm 2.45$ days with $3.3\sigma$ significance. The period uncertainties were estimated using the relation $\Delta P=P\cdot W/T_{\rm span}$, where $P$ is the period in days, $W$ is the duration of the enhanced activity window in days, and $T_{\rm span}=612.33$ days is the total observational time span.
Figure 3C1–C3 present the phase-folded analysis for Period 1. Figure 3C1 shows the intensity profile with both exposure-corrected (green) and uncorrected (blue) rates, along with the Gaussian fit (red dashed). The profile is well-described by a single Gaussian component:
$f_{1}(x)=3951.20\exp\left[-\frac{(x-0.499)^{2}}{2\times(0.0909)^{2}}\right]+867.28.$ (10)
From this Gaussian fit, the active window timing and duration are determined. Based on these results, the green shaded regions in Figure 3A mark the predicted active phases, showing the timing and duration of expected burst windows. Figure 3C2 displays the relative exposure time distribution across phase bins, showing relatively uniform coverage. Figure 3C3 presents the time of arrival (TOA) versus phase diagram, where point sizes are scaled by burst rate weights, demonstrating coherent phase folding over the entire observational baseline.
Figure 3D1–D3 show similar analyses for Period 2. Figure 3D1 reveals a double-peaked structure in the intensity profile, requiring a double-Gaussian model:
@@
[…]
Figure 3D2 shows the exposure time distribution, while Figure 3D3 displays the TOA-phase relation with burst rate weighting.
Notably, Period 2 ($73.60\pm 2.45$ days) is consistent with being the first harmonic of Period 1 ($143.40\pm 7.19$ days), with the ratio $P_{1}/P_{2}=1.95\pm 0.12$ compatible with a 2:1 relationship within uncertainties. The phase-folded diagrams in Figure 3C3 and Figure 3D3 support the interpretation that Period 1 (143.40 days) is the primary period, with Period 2 arising as its harmonic. Therefore, all subsequent discussion will focus solely on Period 1.
@@ 3.1 Method
For periodicity detection of bursts, we adopted the Bayesian framework proposed by the reference [2013ApJ...768...87H]. The method begins by computing the periodogram of burst time series via the FFT, then fits a null hypothesis model to obtain a maximum a posteriori (MAP) parameters. We calculate residuals $R_{j}=2I_{j}/S_{j}$, where $I_{j}$ represents the observed periodogram and $S_{j}$ denotes the null model’s prediction. Markov Chain Monte Carlo (MCMC) sampling generates simulated periodograms, establishing a distribution of maximum residuals against which observed residuals are compared for significance assessment. In order to detect QPOs, we further introduce a nested alternative hypothesis model, computing the likelihood ratio test (LRT) statistic $T_{\mathrm{LRT}}^{\mathrm{obs}}$ and determining its significance through simulated periodograms.
@@
For periodicity detection of burst clusters, we employed the ACF method proposed by the reference [2024NatAs...8..230K]. The method first requires determining an appropriate time window. By analyzing the distribution of burst waiting times, we found that the distribution exhibits a bimodal structure with a transition around 0.2 seconds. Based on this, bursts with intervals shorter than 0.2 seconds are clustered into the same time series. Periodic signals in ACF analysis manifest as equally spaced local maxima. We quantify the strength of periodicity using the statistic $\mathcal{P}=\sum\rho(P_{\mu}\cdot i)$, where $\rho$ denotes the power of the ACF, and $P_{\mu}\cdot i$ represents integer multiples of the period $P_{\mu}$. The original subpulse counts, widths, and amplitudes are preserved, while uniformly distributed perturbations are incorporated at periodic intervals. The simulated sequences then undergo the same ACF analysis to generate a distribution for comparison with the observed values, thereby obtaining the $P$-value.
@@
The periodicity detection method for individual bursts circumvents the challenges of decomposing complex signals into subcomponents. Conversely, the method for burst clusters leverages the timing characteristics of discrete events to facilitate the assessment of periodic significance. For details of the methodology, see references [2022Natur.607..256C, 2024NatAs...8..230K] for the ACF-based analysis and C for the implementation of the FFT-based approach.
Figure 5: Analysis of an FRB burst containing a QPO component using the FFT-based Bayesian inference method. The six panels illustrate key steps in data processing and model evaluation. (a) Frequency–time dynamic spectrum of the burst, dedispersed using a dispersion measure of 528.5 pc cm-3, with RFI-contaminated channels masked; (b) Time series obtained by frequency integration of the dynamic spectrum, used as input for the periodicity analysis; (c) Periodogram of the time series. The blue dashed line represents the background power spectrum modeled under the null hypothesis $H_{0}$ (no QPO component), while the red dashed line corresponds to the alternative hypothesis $H_{1}$ (with a QPO component). A prominent excess of power is seen at a specific frequency. The best-fit parameters for the $H_{1}$ model are listed in Table 2, entry No. 1. (d) Residuals of the periodogram under both $H_{0}$ and $H_{1}$, used to quantify deviations from the respective model fits; (e) Posterior distribution of model parameters obtained via MCMC sampling under $H_{0}$, used to generate simulated periodograms; (f) Distribution of likelihood ratio test statistics computed from the simulated periodograms. The red line marks the observed test statistic $T_{\mathrm{LRT}}^{\mathrm{obs}}$, which quantifies the significance of the QPO component. The observed significance of this QPO, as inferred from panel (f), is $3.4\sigma$.
@@ 3.2 Results
For the ACF analysis, we clustered bursts using a 0.2-second time window, resulting in 516 time series. The ACF was calculated for each of these time series, and the minimum lag of the significant peaks was identified as the candidate period $P_{\mu}$. The significance of these periodicities was assessed using the methodology outlined in Section 3.1. We identified 4 candidates with significance greater than $3\sigma$. These burst time series, their dynamic spectra, ACFs, and distributions of simulated statistics are shown in Figure 4. These time series exhibit relatively few peaks, with those peaks spaced at regular intervals. Most of these candidates exhibit 3 peaks, with one candidate showing 4 peaks, resulting in a high significance of up to $3.9\sigma$. In Figure 7 in D, we also list 7 additional candidates with significance lower than $3\sigma$ yet still showing some evidence of periodicity. Notably, some of these candidates match the period well up to 3 peaks, although the agreement breaks down after the 4th peak, leading to lower significance. This suggests that the candidates with significance greater than $3\sigma$ may not exhibit periodic patterns in longer time series. The periods of these candidates range from 5 ms to 62 ms, reflecting the clustering of bursts within a 0.2-second time window. As expected, increasing the time window length makes it more difficult to maintain periodicity. This aligns with the results from searching for short-timescale periods in the TOA data, where pulse phases exhibited considerable spread rather than remaining confined to a narrow range. The 0.2-second clustering assumption implies that bursts occurring within such short intervals are likely to be physically connected, indicating the potential for periodicity in these local bursts.
Table 2: QPO fitting parameters for individual bursts. The table lists the model components (Model), where components starting with G represent Gaussian models used to fit QPO peaks, and ”Const” denotes the constant model representing the background level. The table also includes their associated amplitudes (Height), central frequencies (CF), full width at half maximum (FWHM), quality factors (Q), constant (C), and significance (Sig.) for the QPO in each individual burst. The MJD in each row indicates the start time of the time series for that particular burst.
No. MJD DM Sig. Model Height CF FWHM Q C ($\times 10^{-3}$)
From 60375 $\rm pc/cm^{3}$ Hz Hz $\times 10^{-3}$
1 0.182604131 528.5 3.4 G0 6.3370 $\pm$ 0.0460 4.45 $\pm$ 0.59 85.02 $\pm$ 0.63 0.0523 $\pm$ 0.0069 -
G1 0.1865 $\pm$ 0.0037 266.85 $\pm$ 0.68 97.00 $\pm$ 1.30 2.7500 $\pm$ 0.0370 -
G2 0.0053 $\pm$ 0.0011 502.10 $\pm$ 6.20 70.00 $\pm$ 14.00 7.1000 $\pm$ 1.4000 -
Const - - - - $3.0\pm 4.6$
2 1.143877879 528.7 3.7 G0 36.3000 $\pm$ 0.4000 5.80 $\pm$ 1.10 118.50 $\pm$ 1.10 0.0487 $\pm$ 0.0094 -
G1 0.9460 $\pm$ 0.0210 327.94 $\pm$ 0.93 117.40 $\pm$ 1.70 2.7930 $\pm$ 0.0410 -
G2 0.0368 $\pm$ 0.0047 579.80 $\pm$ 5.60 105.00 $\pm$ 10.00 5.5000 $\pm$ 0.5500 -
Const - - - - $2.2\pm 1.2$
In the FFT analysis, after excluding radio frequency interference (RFI)—characterized by signals at DM = 0 or their narrowband nature—no periodic signal candidates were found. Nonetheless, two QPOs with significance greater than $3\sigma$ were detected. Figures 5 and Figure 8 display the two bursts containing QPOs. Each figure includes (a) the RFI-masked dynamic spectrum, (b) the dedispersed time series, (c) the periodogram with model fits under the null and alternative hypotheses, (d) residuals from the fits, (e) posterior distributions of model parameters from MCMC sampling, and (f) likelihood ratio test statistic distributions highlighting the observed values. In the periodograms, prominent excess power at specific frequencies indicates the presence of QPO components, with model residuals, MCMC posteriors, and likelihood ratio tests collectively confirming their statistical significance. To assess the statistical significance of the QPOs, we applied the Bayesian method described in Section 3.1 and C. We performed hypothesis testing, where the null hypothesis $H_{0}$ assumed the periodogram could be fitted with one Gaussian model and one constant model, and the apparent QPOs were due to statistical fluctuations. The alternative hypothesis $H_{1}$ assumed that the statistical fluctuations under $H_{0}$ could not explain the apparent QPOs, requiring at least one additional Gaussian model for the QPOs. The LRT statistic yielded observed values of $T_{\rm LRT}^{\rm obs}=34.2$ and $35.0$ for the two QPO events, respectively. Using MCMC sampling, we generated $1\times 10^{6}$ simulated periodograms to compute the simulated likelihood ratios, which were compared with the observed value. As shown in Figure 5f and Figure 8f, the observed likelihood ratios for QPO1 and QPO2 lie in the tail of the simulated distributions under the null hypothesis, indicating that such values are unlikely to arise from statistical fluctuations alone. The corresponding $P$-values are $5.50\times 10^{-4}$ and $1.53\times 10^{-4}$, yielding detection significances of approximately $3.4\sigma$ and $3.7\sigma$, respectively. This result rejects the null hypothesis in favor of the alternative hypothesis, supporting the presence of significant QPOs. Table 2 lists the fitting parameters of periodograms for the QPOs.